# F5-TTS: A Fairytaler that Fakes Fluent and Faithful Speech with Flow Matching

### v2

@@ 6 Conclusion
This work introduces F5-TTS, a fully non-autoregressive text-to-speech system based on flow matching with diffusion transformer (DiT). With a tidy pipeline, literally text in and speech out, F5-TTS achieves state-of-the-art zero-shot ability compared to existing works trained on industry-scale data. We adopt ConvNeXt for text modeling and propose the test-time Sway Sampling strategy to further improve the robustness of speech generation and inference efficiency. Our design allows faster training and inference, by achieving a test-time RTF of 0.15, which is competitive with other heavily optimized TTS models of similar performance. We open-source our code, and models, to enhance transparency and facilitate reproducible research in this area.
Acknowledgments
The authors would like to express our gratitude to Tianrui Wang, Xiaofei Wang, Yakun Song, Yifan Yang, Yiwei Guo, and Yunchong Xiao for the valuable discussions.
@@ Appendix A Baseline Details
VALL-E 2 [13] A large-scale TTS model shares the same architecture as VALL-E [9] but employs a repetition-aware sampling strategy that promotes more deliberate sampling choices, trained on Libriheavy [71] 50K hours English dataset. We compared with results reported in [29].
@@
MELLE [29] An autoregressive large-scale model leverages continuous-valued tokens with variational inference for text-to-speech synthesis. Its variants allow to prediction of multiple mel-spectrogram frames at each time step, noted by MELLE-Rx with x denotes reduction factor. The model is trained on Libriheavy [71] 50K hours English dataset. We compared with results reported in [29].
@@
Voicebox [14] A non-autoregressive large-scale model based on flow matching trained with infilling task. We compared with the 330M parameters trained on 60K hours dataset English-only model’s results reported in [14] and  [12].
@@
NaturalSpeech 3 [12] A non-autoregressive large-scale TTS system leverages a factorized neural codec to decouple speech representations and a factorized diffusion model to generate speech based on disentangled attributes. The 500M base model is trained on Librilight [72] a 60K hours English dataset. We compared with scores reported in [12].
@@
DiTTo-TTS [35] A large-scale non-autoregressive TTS model uses a cross-attention Diffusion Transformer and leverages a pretrained language model to enhance the alignment. We compare with DiTTo-en-XL, a 740M model trained on 55K hours English-only dataset, using scores reported in [35].
@@
FireRedTTS [64] A foundation TTS framework for industry-level generative speech applications. The autoregressive text-to-semantic token model has 400M parameters and the token-to-waveform generation model has about half the parameters. The system is trained with 248K hours of labeled speech data. We use the official code and pre-trained checkpoint to evaluate [^6] https://github.com/FireRedTeam/FireRedTTS.
@@
MaskGCT [66] A large-scale non-autoregressive TTS model without precise alignment information between text and speech following the mask-and-predict learning paradigm. The model is multi-stage, with a 695M text-to-semantic model (T2S) and then a 353M semantic-to-acoustic (S2A) model. The model is trained on Emilia [57] dataset with around 100K Chinese and English in-the-wild speech data. We compare with results reported in [66].
@@
Seed-TTS [39] A family of high-quality versatile speech generation models trained on unknown tremendously large data that is of orders of magnitudes larger than the previously largest TTS systems [39]. Seed-TTSDiT is a large-scale fully non-autoregressive model. We compare with results reported in [39].
@@
E2 TTS [36] A fully non-autoregressive TTS system proposes to model without the phoneme-level alignment in Voicebox, originally trained on Libriheavy [71] 50K English dataset. We compare with our reproduced  [^3] M multilingual E2 TTS trained on Emilia [57] dataset with around 100K Chinese and English in-the-wild speech data.
@@
CosyVoice [65] A two-stage large-scale TTS system, first autoregressive text-to-token, then a flow matching diffusion model. The model is of around 300M parameters, trained on 170K hours of multilingual speech data. We obtain the evaluation result with the official code and pre-trained checkpoint [^7] https://huggingface.co/model-scope/CosyVoice-300M.
@@ B.1 Small Model Configuration
As mentioned in Sec.3.2, F5-TTS leverages an adaLN DiT backbone, while E2 TTS is a flat U-Net equipped Transformer. F5-TTS$+$LongSkip adds an additional long skip structure connecting the first to the last layer in the Transformer. For the Multi-Model Diffusion Transformer (MMDiT) [37], a double stream transformer, the setting denotes one stream configuration.
Table 3: Details of small model configurations.
@@
F5-TTS--Conv2Text 768,18,12,2 - 153M 164
F5-TTS$+$Conv2Audio 768,16,12,2 512,4,2 163M 181
F5-TTS$+$LongSkip 768,18,12,2 512,4,2 159M 175
@@
E2 TTS$+$Conv2Text 768,20,12,4 512,4,2 161M 301
MMDiT [37] 512,16,16,2 - 151M 104
@@ B.2 Ablation study on Input Condition
The ablation study on different input conditions is conducted with three settings: common input with text and audio prompts, providing ground truth duration information rather than an estimate, and retaining only text input dropping audio prompt. In Tab.4, all evaluations take the 155M small models’ checkpoints trained on WenetSpeech4TTS Premium at 800K updates.
Model Common Input Ground Truth Dur. Drop Audio Prompt
@@
F5-TTS$+$Conv2Audio 5.78 0.55 5.28 0.55 3.78 0.21
F5-TTS$+$LongSkip 5.17 0.53 5.03 0.53 3.35 0.21
@@
E2 TTS$+$Conv2Text 18.10 0.49 17.94 0.49 3.06 0.21
B.3 Comparison of ODE Solvers
The comparison results of using the Euler or midpoint ODE solver during F5-TTS inference are shown in Tab.B.3. The Euler is inherently faster (first-order) and performs slightly better typically for larger NFE inference with Sway Sampling (otherwise the Euler solver results in degradation).
Table 5: Evaluation results of F5-TTS (F5) on LibriSpeech-PC test-clean, Seed-TTS test-en and Seed-TTS test-zh, employing the Euler or midpoint ODE solver, and with different Sway Sampling $s$ values. The Real-Time Factor (RTF) is computed with the inference time of 10s speech.
Model   LibriSpeech-PC test-clean Seed-TTS test-en Seed-TTS test-zh
WER(%)↓ SIM-o ↑ WER↓ SIM-o ↑ WER↓ SIM-o ↑ RTF ↓
Ground Truth 2.23 0.69 2.06 0.73 1.26 0.76 -
[…]
F5 (16 NFE Euler) 2.53 0.66 1.89 0.67 1.74 0.75 0.15
F5 (16 NFE midpoint) 2.43 0.66 1.88 0.66 1.61 0.75 0.26
F5 (32 NFE Euler) 2.42 0.66 1.83 0.67 1.56 0.76 0.31
F5 (32 NFE midpoint) 2.41 0.66 1.87 0.66 1.58 0.75 0.53
[…]
F5 (16 NFE Euler) 2.82 0.65 2.14 0.65 2.28 0.72 0.15
F5 (16 NFE midpoint) 2.58 0.65 1.86 0.65 1.70 0.73 0.26
F5 (32 NFE Euler) 2.50 0.66 1.81 0.67 1.62 0.75 0.31
F5 (32 NFE midpoint) 2.42 0.66 1.84 0.66 1.62 0.75 0.53
Table 6: Base model evaluation results on LibriSpeech-PC test-clean, Seed-TTS test-en and test-zh, with and without proposed test-time Sway Sampling (SS, with coefficient $s=-1$) strategy for flow steps. All generations leverage the midpoint ODE solver for ease of ablation.
Model WER(%)↓ SIM-o ↑ RTF ↓
@@ B.3 Comparison of ODE Solvers
Ground Truth (1127 samples) 2.23 0.69 -
Vocoder Resynthesized 2.32 0.66 -
[…]
E2 TTS (32 NFE w/ SS) 2.84 0.72 0.68
E2 TTS (32 NFE w/o SS) 2.95 0.69 0.68
[…]
F5-TTS (32 NFE w/ SS) 2.41 0.66 0.53
F5-TTS (32 NFE w/o SS) 2.84 0.62 0.53
@@
Ground Truth (1088 samples) 2.06 0.73 -
Vocoder Resynthesized 2.09 0.70 -
[…]
E2 TTS (32 NFE w/ SS) 1.98 0.73 0.68
E2 TTS (32 NFE w/o SS) 2.19 0.71 0.68
[…]
F5-TTS (32 NFE w/ SS) 1.87 0.66 0.53
F5-TTS (32 NFE w/o SS) 1.93 0.63 0.53
@@
Ground Truth (2020 samples) 1.26 0.76 -
Vocoder Resynthesized 1.27 0.72 -
[…]
E2 TTS (32 NFE w/ SS) 1.77 0.78 0.68
E2 TTS (32 NFE w/o SS) 1.97 0.73 0.68
[…]
F5-TTS (32 NFE w/ SS) 1.58 0.75 0.53
F5-TTS (32 NFE w/o SS) 1.93 0.69 0.53
B.4 Sway Sampling Effectiveness on Base Models
From Tab.B.3, it is clear that our Sway Sampling strategy for test-time flow steps consistently improves the zero-shot generation performance in aspects of faithfulness to prompt text (WER) and speaker similarity (SIM). The gain of applying Sway Sampling to E2 TTS [36] proves that our Sway Sampling strategy is universally applicable to existing flow matching based TTS models.
@@ B.5 ELLA-V Hard Sentences Evaluation
ELLA-V [15] proposed a challenging set containing 100 difficult textual patterns evaluating the robustness of the TTS model. Following previous works [13, 29, 36], we include generated samples in our demo page [^8] https://SWivid.github.io/F5-TTS. We additionally compare our model with the objective evaluation results reported in E1 TTS [49].

### v3

@@ 6 Conclusion
This work introduces F5-TTS, a fully non-autoregressive text-to-speech system based on flow matching with diffusion transformer (DiT). With a tidy pipeline, literally text in and speech out, F5-TTS achieves state-of-the-art zero-shot ability compared to existing works trained on industry-scale data. We adopt ConvNeXt for text modeling and propose the test-time Sway Sampling strategy to further improve the robustness of speech generation and inference efficiency. Our design allows faster training and inference, by achieving a test-time RTF of 0.15, which is competitive with other heavily optimized TTS models of similar performance. We will open-source our code, and models, to enhance transparency and facilitate reproducible research in this area.
Limitations
There are two limitations to this work. First, although F5-TTS accelerates the training and inference while maintaining the simplicity of the system through better structure design, the mel spectrogram sequence length is still much longer than the text modality. Therefore, research and employment of a more efficient and hopefully universal continuous representation compatible with highly expressive speech synthesis remains a critical direction and can further improve efficiency and performance. Second, although F5-TTS has great zero-shot generation ability and can deeply mimic the reference audio, it lacks fine-grained control of paralinguistic details, e.g. emotion, which is of great research and practical application value.
@@ Appendix A Baseline Details
VALL-E 2 (Chen et al., 2024) A large-scale TTS model shares the same architecture as VALL-E (Wang et al., 2023a) but employs a repetition-aware sampling strategy that promotes more deliberate sampling choices, trained on Libriheavy (Kang et al., 2024) 50K hours English dataset. We compared with results reported in Meng et al. (2024).
@@
MELLE (Meng et al., 2024) An autoregressive large-scale model leverages continuous-valued tokens with variational inference for text-to-speech synthesis. Its variants allow to prediction of multiple mel-spectrogram frames at each time step, noted by MELLE-Rx with x denotes reduction factor. The model is trained on Libriheavy (Kang et al., 2024) 50K hours English dataset. We compared with results reported in Meng et al. (2024).
@@
Voicebox (Le et al., 2024) A non-autoregressive large-scale model based on flow matching trained with infilling task. We compared with the 330M parameters trained on 60K hours dataset English-only model’s results reported in Le et al. (2024) and  Ju et al. (2024).
@@
NaturalSpeech 3 (Ju et al., 2024) A non-autoregressive large-scale TTS system leverages a factorized neural codec to decouple speech representations and a factorized diffusion model to generate speech based on disentangled attributes. The 500M base model is trained on Librilight (Kahn et al., 2020) a 60K hours English dataset. We compared with scores reported in Ju et al. (2024).
@@
DiTTo-TTS (Lee et al., 2024) A large-scale non-autoregressive TTS model uses a cross-attention Diffusion Transformer and leverages a pretrained language model to enhance the alignment. We compare with DiTTo-en-XL, a 740M model trained on 55K hours English-only dataset, using scores reported in Lee et al. (2024).
@@
FireRedTTS (Guo et al., 2024a) A foundation TTS framework for industry-level generative speech applications. The autoregressive text-to-semantic token model has 400M parameters and the token-to-waveform generation model has about half the parameters. The system is trained with 248K hours of labeled speech data. We use the official code and pre-trained checkpoint to evaluate [^5] https://github.com/FireRedTeam/FireRedTTS.
@@
MaskGCT (Wang et al., 2024) A large-scale non-autoregressive TTS model without precise alignment information between text and speech following the mask-and-predict learning paradigm. The model is multi-stage, with a 695M text-to-semantic model (T2S) and then a 353M semantic-to-acoustic (S2A) model. The model is trained on Emilia (He et al., 2024) dataset with around 100K Chinese and English in-the-wild speech data. We compare with results reported in Wang et al. (2024).
@@
Seed-TTS (Anastassiou et al., 2024) A family of high-quality versatile speech generation models trained on unknown tremendously large data that is of orders of magnitudes larger than the previously largest TTS systems (Anastassiou et al., 2024). Seed-TTSDiT is a large-scale fully non-autoregressive model. We compare with results reported in Anastassiou et al. (2024).
@@
E2 TTS (Eskimez et al., 2024) A fully non-autoregressive TTS system proposes to model without the phoneme-level alignment in Voicebox, originally trained on Libriheavy (Kang et al., 2024) 50K English dataset. We compare with our reproduced  [^3] M multilingual E2 TTS trained on Emilia (He et al., 2024) dataset with around 100K Chinese and English in-the-wild speech data.
@@
CosyVoice (Du et al., 2024b) A two-stage large-scale TTS system, first autoregressive text-to-token, then a flow matching diffusion model. The model is of around 300M parameters, trained on 170K hours of multilingual speech data. We obtain the evaluation result with the official code and pre-trained checkpoint [^6] https://huggingface.co/model-scope/CosyVoice-300M.
@@ Appendix B Experimental Result Supplements
The UTMOS (Saeki et al., 2022) scores reported in this section are evaluated with an open-source MOS prediction model [^7] https://github.com/tarepan/SpeechMOS. The UTMOS is an objective metric measuring naturalness.
@@ B.1 Small Model Configuration
As mentioned in Sec.3.2, F5-TTS leverages an adaLN DiT backbone with ConvNeXt V2 blocks, while E2 TTS is a flat U-Net equipped Transformer. F5-TTS++LongSkip adds an additional long skip structure connecting the first to the last layer (Lee et al., 2024) in the Transformer. For the Multi-Model Diffusion Transformer (MMDiT) (Esser et al., 2024), a double stream transformer, the setting denotes one stream configuration.
@@
F5-TTS-Conv2Text    768,18,12,2 - 153M 164
F5-TTS++Conv2Audio  768,16,12,2 512,4,2 163M 181
F5-TTS++LongSkip    768,18,12,2 512,4,2 159M 175
@@
E2 TTS++Conv2Text    768,20,12,4 512,4,2 161M 301
MMDiT        512,16,16,2 - 151M 104
Table 3: Details of small model configurations.
@@ B.2 Ablation study on Input Condition
The ablation study on different input conditions is conducted with three settings:
  • Common input with text and audio prompts.
  • Providing ground truth duration information rather than an estimate.
  • Retaining only text input, dropping audio prompt (using blank).
In Tab.4, all evaluations take the 155M small models’ checkpoints trained on WenetSpeech4TTS Premium at 800K updates. Analysis see Sec.5.1.
Model        Common Input   GT Duration   Text-Only
@@
F5-TTS++Conv2Audio  5.78 0.55 5.28 0.55 3.78 0.21
F5-TTS++LongSkip    5.17 0.53 5.03 0.53 3.35 0.21
@@
E2 TTS++Conv2Text   18.10 0.49 17.94 0.49 3.06 0.21
B.3 Sway Sampling Effectiveness on Base Models
From Tab.5, it is clear that our Sway Sampling strategy for test-time flow steps consistently improves the zero-shot generation performance in aspects of faithfulness to prompt text (WER), speaker similarity (SIM), and naturalness (UTMOS). The gain of applying Sway Sampling to E2 TTS (Eskimez et al., 2024) proves that our Sway Sampling strategy is universally applicable to existing flow matching based TTS models.
Model         WER(%)↓ SIM-o↑ UTMOS↑ RTF↓
@@ B.3 Sway Sampling Effectiveness on Base Models
Ground Truth 2.23 0.69 4.09 -
Vocoder Resynthesized 2.32 0.66 3.64 -
E2 TTS (16 NFE w/ SS) 2.86 0.71 3.66 0.34
E2 TTS (32 NFE w/ SS) 2.84 0.72 3.70 0.68
E2 TTS (32 NFE w/o SS) 2.95 0.69 3.56 0.68
F5-TTS (16 NFE w/ SS) 2.43 0.66 3.87 0.26
F5-TTS (32 NFE w/ SS) 2.41 0.66 3.89 0.53
F5-TTS (32 NFE w/o SS) 2.84 0.62 3.70 0.53
@@
Ground Truth 2.06 0.73 3.53 -
Vocoder Resynthesized 2.09 0.70 3.33 -
E2 TTS (16 NFE w/ SS) 1.99 0.72 3.55 0.34
E2 TTS (32 NFE w/ SS) 1.98 0.73 3.57 0.68
E2 TTS (32 NFE w/o SS) 2.19 0.71 3.33 0.68
F5-TTS (16 NFE w/ SS) 1.88 0.66 3.70 0.26
F5-TTS (32 NFE w/ SS) 1.87 0.66 3.72 0.53
F5-TTS (32 NFE w/o SS) 1.93 0.63 3.51 0.53
@@
Ground Truth 1.26 0.76 2.78 -
Vocoder Resynthesized 1.27 0.72 2.61 -
E2 TTS (16 NFE w/ SS) 1.80 0.78 2.84 0.34
E2 TTS (32 NFE w/ SS) 1.77 0.78 2.87 0.68
E2 TTS (32 NFE w/o SS) 1.97 0.73 2.49 0.68
F5-TTS (16 NFE w/ SS) 1.61 0.75 2.87 0.26
F5-TTS (32 NFE w/ SS) 1.58 0.75 2.91 0.53
F5-TTS (32 NFE w/o SS) 1.93 0.69 2.58 0.53
Table 5: Base model evaluation results on LibriSpeech-PC test-clean, Seed-TTS test-en and test-zh, with and without proposed test-time Sway Sampling ($S$, with coefficient s=1s=-1) strategy for flow steps. All generations leverage the midpoint ODE solver for ease of ablation.
B.4 Comparison of ODE Solvers
The comparison results of using the Euler (first-order), midpoint (second-order), or improved Heun (third-order, Heun-3) ODE solver during F5-TTS inference are shown in Tab.6. The Euler is inherently faster and performs slightly better typically for larger NFE inference with Sway Sampling (otherwise the Euler solver results in degradation).
LibriSpeech-PC test-clean Seed-TTS test-en Seed-TTS test-zh
F5-TTS WER(%)↓ SIM-o↑ UTMOS↑ WER(%)↓ SIM-o↑ UTMOS↑ WER(%)↓ SIM-o↑ UTMOS↑ RTF↓
Ground Truth 2.23 0.69 4.09 2.06 0.73 3.53 1.26 0.76 2.78 -
s=1s=-1
16 NFE Euler 2.53 0.66 3.88 1.89 0.67 3.76 1.74 0.75 2.96 0.15
16 NFE midpoint 2.43 0.66 3.87 1.88 0.66 3.70 1.61 0.75 2.87 0.26
32 NFE Euler 2.42 0.66 3.90 1.83 0.67 3.76 1.56 0.76 2.95 0.31
32 NFE midpoint 2.41 0.66 3.89 1.87 0.66 3.72 1.58 0.75 2.91 0.53
16 NFE Heun-3 2.39 0.65 3.87 1.80 0.66 3.70 1.55 0.75 2.88 0.44
s=0.8s=-0.8
16 NFE Euler 2.82 0.65 3.73 2.14 0.65 3.70 2.28 0.72 2.74 0.15
16 NFE midpoint 2.58 0.65 3.86 1.86 0.65 3.68 1.70 0.73 2.83 0.26
32 NFE Euler 2.50 0.66 3.89 1.81 0.67 3.74 1.62 0.75 2.94 0.31
32 NFE midpoint 2.42 0.66 3.89 1.84 0.66 3.70 1.62 0.75 2.91 0.53
16 NFE Heun-3 2.40 0.65 3.85 1.78 0.66 3.68 1.56 0.74 2.84 0.44
Table 6: Evaluation results of F5-TTS on LibriSpeech-PC test-clean, Seed-TTS test-en and Seed-TTS test-zh, employing the Euler, midpoint or Heun-3 ODE solver, and with different Sway Sampling $s$ values.
@@ B.5 ELLA-V Hard Sentences Evaluation
ELLA-V (Song et al., 2024) proposed a challenging set containing 100 difficult textual patterns evaluating the robustness of the TTS model. Following previous works (Chen et al., 2024; Meng et al., 2024; Eskimez et al., 2024), we include generated samples in our demo page [^8] https://SWivid.github.io/F5-TTS/. We additionally compare our model with the objective evaluation results reported in E1 TTS (Liu et al., 2024b).
@@
B.6 Comparison of Vocoders and between PC and non-PC
The inference results with pretrained BigVGAN (Lee et al., 2022) and Vocos (Siuzdak, 2023) respectively as vocoder are shown in Tab.8, along with additional evaluation on a non-Capitalized version removing all Punctuations (non-PC) of the filtered LibriSpeech-PC test-clean subset. The non-PC version equals an ordinary LibriSpeech test-clean subset, with which we provide more comprehensive comparisons with previous works.
Moreover, we include WER scores measuring with a Hubert-large-based (Hsu et al., 2021) ASR model [^9] https://huggingface.co/facebook/hubert-large-ls960-ft in Tab.8, with which our reproduced multilingual E2 TTS with 32 NFE and Vocos as vocoder achieves a WER of 2.92 on LibriSpeech-PC test-clean and 2.66 if Sway Sampling applied.
NFE steps LibriSpeech-PC test-clean LibriSpeech-non-PC test-clean Seed-TTS test-en Seed-TTS test-zh
& Vocoder WER(%)↓ SIM-o ↑ WER(%)↓ SIM-o ↑ WER(%)↓ SIM-o ↑ WER(%)↓ SIM-o ↑
Ground Truth 2.23 (1.89) 0.69 2.29 (1.86) 0.69 2.06 0.73 1.26 0.76
16 NFE - Vocos 2.53 (2.34) 0.66 2.72 (2.53) 0.66 1.89 0.67 1.74 0.75
16 NFE - BigVGAN 2.21 (1.96) 0.67 2.55 (2.34) 0.67 1.65 0.66 1.64 0.74
32 NFE - Vocos 2.42 (2.09) 0.66 2.44 (2.16) 0.66 1.83 0.67 1.56 0.76
32 NFE - BigVGAN 2.11 (1.81) 0.67 2.28 (2.03) 0.67 1.62 0.66 1.53 0.74
Table 8: F5-TTS Base model evaluation results on LibriSpeech-PC test-clean, LibriSpeech-non-PC test-clean, Seed-TTS test-en and test-zh with BigVGAN and Vocos, default setting as in Sec.5. The WER scores in brackets indicate results leveraging the Hubert-large-based ASR model.
Train Set LibriTTS - 585 hours LJSpeech - 24 hours
Test Set LibriSpeech-PC test-clean LJSpeech in-set tests
Model (#Param.) Update WER(%)↓ SIM-o ↑ UTMOS ↑ WER(%)↓ SIM-o ↑ UTMOS ↑
Ground Truth - 2.23 0.69 4.09 2.36 0.72 4.36
USLM (361M) - 6.1 0.43 - - - -
100K 29.5 0.53 3.78 5.64 0.72 4.17
200K 4.58 0.59 4.07 2.93 0.72 4.18
F5-TTS small 300K 2.71 0.60 4.11 3.26 0.71 4.12
(158M) 400K 2.44 0.60 4.11 3.90 0.70 4.05
500K 2.20 0.60 4.10 4.68 0.70 3.99
600K 2.23 0.59 4.10 5.25 0.69 3.93
Table 9: F5-TTS small models evaluation results on LibriSpeech-PC test-clean (model trained on LibriTTS 585 hours multi-speaker dataset), and on LJSpeech in-set test samples (model trained on 24 hours single-speaker LJSpeech); Vocos as vocoder, Whisper-large-v3 as ASR model. The scores of USLM (Zhang et al., 2023a) are evaluated with the official checkpoint pre-trained on LibriTTS.
B.7 Training and Inference Performance with Different Dataset Scales
We train F5-TTS 158M small models on LibriTTS (Zen et al., 2019) 585 hours and LJSpeech (Ito and Johnson, 2017) 24 hours English datasets to provide insights on our model’s training stability with different dataset scales, typically to see whether it can maintain stable training on limited data. Both training takes place with the same configuration as described in Sec.5.1 and Appendix B.1 despite a batch size of 307,200 audio frames (0.91 hours) as base models. Every 100K update takes approximately 8 hours on 8 NVIDIA H100 SXM GPUs.
Same as Sec.5, we report the average score of three random seed generation results, using a CFG strength of 2, a Sway Sampling coefficient of 1-1, and 32 NFE steps. Since LJSpeech is a single-speaker dataset, we measure the metrics on in-set tests (1000 samples organized with 4 to 7 seconds to infer with an around 3-second prompt). It is clear from Tab.9 (and Fig.2 in comparison with E2 TTS small) that our design enables stable training to learn speech-text alignment (without grapheme-to-phoneme) with varying data amounts.
Appendix C Subjective Evaluation Details
To evaluate speech quality, we conduct a CMOS subjective evaluation. 20 natives were invited for both English and Mandarin to evaluate 30 rounds with randomly selected utterances for all three test sets and all model variants. Evaluators were informed in detail about the guidelines and scoring criteria for the CMOS test, for example, the general instruction part:
  • Most important: use high-quality studio headphones and a good sound card!
  • Listen through all test files and test sets before you do any ratings to get used to the material.
  • Rate the quality of the test items only compared to the reference on top.
  • Try to rate the overall impression of a test item and don’t concentrate on single aspects.
For the CMOS test, the overall quality of a generated speech is first rated from 3-3 (bad quality compared to the reference) to +3+3 (much better than the reference) integer scale, then reported in average differentials with received scores of ground truth speech. For SMOS, a 1 to 5 with 0.5 as an interval rating is employed (higher better). Judges are to score the similarity between the synthesized and prompt speech with clearly differentiated instructions mentioning:
  • Try to rate concentrating on the speaker similarity aspects with reference speech.
We encourage more rigorous and transparent subjective evaluations, such as releasing used samples if not open-sourcing the model checkpoints. Meanwhile inviting more evaluators leads to more comprehensive and fair rating scores. Just for reference, DiTTo-TTS (Lee et al., 2024) received and reported 6 and 12 ratings for SMOS and CMOS, respectively, NaturalSpeech 3 (Ju et al., 2024) invited 12 natives to judge 20 samples for CMOS and 10 samples for SMOS.